OpenStudy (anonymous):

The sum of two consecutive even integers is 42. Find the numbers.

pooja195 (pooja195):

The two even integers are 20 and 22. 20 + 22 = 42.

pooja195 (pooja195):

ALSO WELCOME TO OPENSTUDY!!! :)

pooja195 (pooja195):

http://openstudy.com/code-of-conduct

terenzreignz (terenzreignz):

Are you wondering at how to arrive at 20 and 22?

pooja195 (pooja195):

well guess n check u kno 20+20 is 40 so is u do 22 u get 42

terenzreignz (terenzreignz):

Proper way to do is ... Let x and x+2 be the even integers. Then x + x + 2 = 42 2x + 2 = 42 2x = 40 x = 20 x+2 = 22